# Dual-Personalizing Adapter for Federated Foundation Models

### v1

@@ 3.2 Test-time Personalization
Let us clarify the concepts of “local dataset”, “personalization” and “test-time”. In our setting, the “local dataset” refers to the data used in the training phase, serving as the source domain. “Personalization” denotes optimal performance in this source domain. The term “test-time” refers to the presence of additional data characterized by distribution shifts, used exclusively for testing. In our setting, test-time personalization aims to achieve the personalization while ensuring comparable performance on tasks encountered during the test-time phase.
4 Proposed Method
@@
To simplify the illustration, we use LLM as the backbone of our proposed framework and adopt LoRA Hu et al. (2021) as the adapter-based PEFT method in our framework. The overall framework is easy to adapt to other types of backbone and other adapter-based PEFT methods. LoRA decomposes the training weight into a frozen weight ${\bm{\theta}}$, and a trainable weight derived through the multiplication of two low-rank weights $\Delta{\bm{\theta}}=\Delta{\bm{\theta}}^{b}\Delta{\bm{\theta}}^{a}$. Given the inherent data heterogeneity intrinsic to federated learning environments, distinct NLP tasks are allocated to different clients. Consequently, within our FL framework, this data heterogeneity is manifest primarily in the distribution characteristics of the inputs across these diverse NLP tasks. Next, we will delve deeply into our proposed framework.
4.1 Overall Framework
In order to align with the application scenarios, we consider the test-time personalization setting, where the test-time distribution shifts data are tested. To simplify, we consider test-time personalization in an ideal setting, where all tasks are included in all clients. Therefore, for each client, the local task is taken as the primary task for personalization, while tasks from other clients are taken as the test-time tasks during the testing phase.
This framework raises two pivotal considerations: personalization and test-time distribution shifts. Since an aggregated global model at the server may not adequately address the specific tasks of each client, we introduce a local model to achieve personalization, with the premise that this personalization is anchored in maintaining local capabilities through training on each client’s local dataset. However, the aspect of distribution shifts during the testing phase remains a challenge, as conventional personalization methods often falter in these test-time tasks. Considering that data from other clients represents these test-time tasks, we can glean insights from the aggregated global model, which has already acquired a comprehensive understanding of all tasks through aggregation algorithms.
Therefore, we propose a dual-personalizing adapter (FedDPA) system for each client, based on the intuition that a global adapter targeting the test-time tasks and a local adapter for personalization are tuned together under a sophisticated federated learning algorithm. Additionally, we investigate two personalization methods to learn the local adapter, dubbed FedDPA-F and FedDPA-T respectively. As illustrated in Fig 1, the global adapter is obtained by conventional FL training and the local adapter is maintained locally by tuning on each local dataset. These two components are strategically combined to realize test-time personalization.
4.2 FL Training of Global Model
@@
Addressing test-time distribution shifts necessitates a comprehensive acquisition of general knowledge. The conventional Federated Learning process is inherently designed to aggregate this general knowledge from a diverse range of tasks. Consequently, we utilize the adapter trained within the FL context as the global adapter for addressing test-time tasks, with more comprehensive details provided subsequently.
@@
As illustrated in Fig 1, during the training stage, each client is endowed with a unique task and its corresponding local dataset ${\mathbb{D}}^{m}$, which is harnessed for model training. At each client, there consists of a frozen LLM model $f_{m}({\bm{x}};{\bm{\theta}})$ with a global lightweight global adapter (LoRA) $\Delta{\bm{\theta}}_{m}=\Delta{\bm{\theta}}^{b}_{m}\Delta{\bm{\theta}}^{a}_{m}$. This global adapter is used for aggregation by sending to the server. Notably, the server’s role is limited to computing the aggregated adapter $\Delta{\bm{\theta}}$, thus obviating the need for maintaining a large-scale model. During each communication round $k$, the global adapter $\Delta{\bm{\theta}}_{m}^{k}$ is seeded with the aggregated adapter $\Delta{\bm{\theta}}^{k}$ from the central server, denoted as $\Delta{\bm{\theta}}_{m}^{k}=\Delta{\bm{\theta}}^{k}$. After this initialization, the model is trained on local datasets for a specified number of epochs, after which the updated adapter parameters are transmitted back to the server. Upon receipt of the adapter weights $\Delta{\bm{\theta}}_{m}^{k}$ of all activated clients, the server employs FedAvg to aggregate these adapter weights. This process results in the formulation of the updated adapter weights $\Delta{\bm{\theta}}^{k+1}$ for the next round. This iterative paradigm persists until training reaches its convergence. The overall objective can be formulated as:
$\min_{\Delta{\bm{\theta}}}\sum^{M}_{m=1}\sum_{(x,y)\in{\mathbb{D}}^{m}}\frac{1}{M}\mathcal{L}(x,y;{\bm{\theta}};\Delta{\bm{\theta}})$ (3)
@@
Other federated algorithms like FedProx Li et al. (2020) can also be applied with LoRA tuning of this global model learning. In this paper, we just take FedAVG as an example.
@@ 4.3 Personalization of Local Model
The primary purpose of test-time personalization lies in the concept of personalization. To actualize this, we integrate a local adapter into our approach. Drawing inspiration from prior research Tan et al. (2022), We introduce two methods for personalizing the local adapter: 1) directly fine-tuning the local adapter initialized by the global adapter, and 2) learning an additional local adapter during the training. In the first method, the local adapter is initialized by the global adapter, followed by subsequent fine-tuning using local datasets to facilitate personalization. The second method involves maintaining a separate local adapter locally during training without communication. At each stage of communication, this local adapter is initialized based on its previous state and then fine-tuned in conjunction with a frozen global adapter using local datasets, thereby facilitating the acquisition of personalized knowledge.
@@
Figure 2: Frameworks of two personalized methods for the local adapter (LoRA).
@@
To be more specific, a local adapter (LoRA) $\Delta\hat{{\bm{\theta}}}=\Delta\hat{{\bm{\theta}}}^{b}\Delta\hat{{\bm{\theta}}}^{a}$ is introduced. Thus, this model contains three components: a frozen LLM ${\bm{\theta}}$, a global adapter (LoRA) $\Delta{\bm{\theta}}$ and a local adapter (LoRA) $\Delta\hat{{\bm{\theta}}}$. As delineated in Fig 2 (a), for the first method, after training, the local adapter is first initialized by the global adapter denoted as $\Delta\hat{{\bm{\theta}}}=\Delta{\bm{\theta}}$, then fine-tuned on local data to get the final local adapter. As shown in Fig 2 (b), for the second method, during each communication round in training for each adapter layer, upon receiving an input ${\bm{h}}$, it simultaneously traverses the frozen LLM, the frozen global adapter and the local adapter. The process entails an initial fusion of the outputs from both the local and global adapters, followed by integration with the output of the LLM to yield the final result ${\bm{h}}^{{}^{\prime}}$. A predefined weighting factor of $\alpha$ is employed to balance the contributions between the local and global adapters. The computational workflow of this configuration is mathematically formulated as follows:
${\bm{h}}^{{}^{\prime}}={\bm{\theta}}{\bm{h}}+((1-\alpha)\cdot\Delta{\bm{\theta}}{\bm{h}}+\alpha\cdot\Delta\hat{{\bm{\theta}}}{\bm{h}})$ (4)
4.4 Instance-wise Dynamic Weighting Mechanism
Upon training completion, we obtain a global adapter for test-time tasks and a local adapter for the personalized targeted task. The next question becomes how to combine them for subsequent inference tasks. This challenge is particularly pronounced in practical applications where the nature of the input instance—whether it pertains to the targeted task or test-time tasks—is not predetermined. Therefore, dynamically adjusting the weighting ratio between the global and local adapters for each instance becomes necessary. Considering the disparate data distributions that characterize test-time tasks and local tasks, coupled with the wealth of training instances specific to local tasks available to each client, we advocate for a novel approach to calculate the similarity between each incoming instance and the available training instances. This similarity metric serves as the key factor in determining the proportional contribution of the global and local adapters, thereby effectively guiding the combined weighting process and ensuring optimal performance across varying task scenarios.
@@ 4.4 Instance-wise Dynamic Weighting Mechanism
Our proposed instance-wise dynamic weighting mechanism calculates the similarity between the input instance and local instances, using this metric to determine the appropriate weight balance for the global and local adapter combination. To facilitate this, the representation of each input instance is essential. Leveraging the robust capability of pre-trained LLMs to abstract input sentences, we utilize the hidden states from the final layer of the LLM as the representation for similarity computation. Given that the LLM is decoder-based, with tokens attending only to preceding tokens, the embedding of the final token is considered representative of the entire input for similarity evaluation. Furthermore, to enhance the representation quality, the global adapter, which embodies general knowledge, is incorporated into this embedding process.
@@
More specifically, during the inference stage, for each input instance $x$ in a client, we randomly sample $S$ instances $\{x_{0},x_{1},...,x_{s}\}$ from the local training dataset. These instances are then fed into the LLM, augmented with the global adapter, to obtain the last token’s embeddings from the final layer, denoted as ${\bm{w}}_{x}$ and $\{{\bm{w}}_{x_{0}},{\bm{w}}_{x_{1}},...,{\bm{w}}_{x_{s}}\}$ respectively. Subsequently, we calculate the cosine similarity between the input representation ${\bm{w}}_{x}$ and each sampled local representation in $\{{\bm{w}}_{x_{0}},{\bm{w}}_{x_{1}},...,{\bm{w}}_{x_{s}}\}$, resulting in a score range of $[0,1]$. Finally, we average all scores to obtain the final result, represented as $\alpha=\sum_{i=0}^{S}\frac{1}{S}\cos({\bm{w}}_{x},{\bm{w}}_{x_{i}})$.
@@
Through this methodology, the balancing of weights between the global and local adapters is dynamically adjusted for each test instance, ensuring the model not only tailors to the individual client’s specific needs but also benefits from the aggregated model’s collective knowledge across test-time tasks.
@@
Table 1: Personalization and test-time personalization results of different models on federated dataset 1. FedDPA-F represents the model with the local fine-tuning adapter and FedDPA-T represents the model with the local training adapter. Linguistic represents the linguistic acceptability task, Word Dis represents the word disambiguation task, and Question CLS represents the question classification task.
5 Experiment
5.1 Datasets and Baselines
In our experiment, we construct two federated datasets from Flan Wei et al. (2021). Flan is a collection of diverse NLP tasks from over 60 datasets. For each dataset, different templates are used to transfer each example into an instruction for generative turning. In order to be better suitable for FL settings, we randomly select 8 NLP tasks from different datasets for each federated dataset and randomly select 300 examples for training and 200 examples for testing. ROGUE-1 is taken as a metric.
Here, we compare our methods with four baselines based on the same model architecture: centralized model, Local-finetuned model, FedIT Zhang et al. (2023a) and FedLoRA model Yi et al. (2023). The centralized model is trained on all data of tasks in one center. The local-finetuned model infers that only local data are used to train the model without any communication with other clients or the server. Here, we adapt the training paradigm in FedLoRA Yi et al. (2023) to NLP tasks.
5.2 Implementation Details
We distribute the data between clients based on the NLP task for data heterogeneity. Since we select 8 NLP tasks, corresponding to $M=8$ clients in our experiment. To better evaluate the effectiveness of our method, we assume that all clients are activated for every communication round and set the communication round $K$ to 20. For local training, each client conducts 10 local epochs with a batch size of 32. During training, we adapt alpaca-LoRA as our base model and initialize it with LLaMA-7B. The rank of LoRA is set as $r=8$ and only applied to ${\bm{W}}_{q}$ and ${\bm{W}}_{v}$. The updating weight of each client’s local LoRA during training is $\alpha=0.5$ for federated dataset 1 and $\alpha=0.3$ for federated dataset 2. We set $S=5$ to select instances for the Instance-wise dynamic weighting mechanism.
@@ 5.3 Main Results
In this section, we represent the experiments of our methods compared with other baselines. This evaluation encompasses two primary facets: personalization (scores on targeted local tasks) and test-time personalization (average scores on all tasks including targeted local tasks and test-time tasks).
As evidenced in Table 1 and Table 2, our proposed dual-personalizing adapter methods (both fine-tuning and training) exhibit superior performance in personalization compared to other baseline models, which demonstrates the effectiveness of local adapter maintenance for enhancing performance on the targeted local task. For test-time personalization, the FedDPA-F method stands out as the most effective among all personalized models, which suggests that incorporating learning from the global adapter can be instrumental in adapting to test-time distribution shifts for a more comprehensive model achievement. In addition, it is noteworthy that while centralized or global models may yield higher average performances across all tasks, they fall short in excelling at specific tasks for personalization, aligning with the conclusions of the previous study Wang et al. (2023).
@@ 6.1 Convergence Analysis
We present the convergence analysis of our methods in Figure 3. As illustrated in Figure 3 (a), we compared our methods with FedIT and FedLoRA for personalization, with the results showcasing the average performance on target local tasks across all clients. Notably, our methods exhibit a more rapid convergence compared to FedIT and achieve notable performance enhancements after five communication rounds. Despite sharing similar trends with FedLoRA, our approaches, particularly the FedDPA-T, ultimately outperform in personalization. For a more granular insight into test-time personalization convergence, we contrast our approaches with other baselines, delineating average performance on all tasks, including each client’s targeted local and test-time tasks. Figure 3 (b) substantiates that our approaches demonstrate faster convergence rates, further bolstering the efficacy of our methods. Our methods demonstrate parallel trends with FedIT due to the benefit of general knowledge from the global adapter for test-time personalization.
Figure 3: Average accuracy varies as communication rounds. TTP represents test-time personalization.
6.2 Ablation Study
Impact of Instance-Wise Dynamic Weighting Mechanism.
To explore the impact of the instance-wise dynamic weighting mechanism, we implemented experiments with FedDPA methods on different datasets. As shown in Table 3, the incorporation of an instance-wise dynamic weighting mechanism contributes significantly to enhancing performance in both personalization and test-time personalization scenarios. This enhancement is particularly pronounced for FedDPA-T.
@@ 6.2 Ablation Study
Impact of Updating Weight $\alpha$.
In this study, we investigated the influence of the updating weight $\alpha$ during FedDPA-T training with its value $\alpha\in\{03,0.5,0.7\}$. As can be seen in Table 4, for test-time personalization, increasing updating weight $\alpha$ will decrease the performance due to the increased proportion of the local adapter in the model, while for personalization, different updating weights $\alpha$ are required for different datasets to achieve their optimal results.
@@
Impact of Client Sample Rate.
To explore how the number of participated clients impacts model performance, we implemented experiments with sample rate $\{0.2,0.4,0.6,0.8,1\}$. More specifically, we split the data of federated dataset 1 into 5 subsets for each task, where each subset has an equal number of training data and is assigned to one client. For each communication round, the server will select clients from each task based on the sample rate. As shown in Figure 4, as the client participant rates increase, model accuracy also increases as more participating clients provide more data for knowledge learning. Besides, FedDPA-F performs better than FedDPA-T due to the possibility of overfitting when handling a small dataset,
Figure 4: Average accuracy varies as different client participation fractions. TTP represents test-time personalization.

### v2

@@ 4 Proposed Method
To simplify the illustration, we use LLM as the backbone and adopt LoRA hu2021lora as the adapter-based PEFT method in our framework. The overall framework is easy to adapt to other types of backbone and other adapter-based PEFT methods. LoRA decomposes the training weight into a frozen weight ${\bm{\theta}}$, and a trainable weight derived by the multiplication of two low-rank weights $\Delta{\bm{\theta}}=\Delta{\bm{\theta}}^{b}\Delta{\bm{\theta}}^{a}$. The data heterogeneity in FL with LLM primarily manifests as distribution shifts across various NLP tasks among different clients, driven by diverse backgrounds, topics, and other contextual factors, and local loss $\mathcal{L}_{m}$ for all NLP tasks is a standard language modeling objective brown2020language .
4.1 Generic Learning of Global Model
@@
Addressing test-time distribution shifts requires the acquisition of generic knowledge that is applicable across various distributions arjovsky2020out . The conventional federated learning process is inherently designed to aggregate this generic knowledge among different non-IID data. Consequently, we utilize the adapter trained within the FL context as the global adapter for addressing test-time tasks. To further enhance generic learning, our model aggregation strategy is based on the client number rather than the number of data by considering the potential biases stemming from different numbers of tasks.
@@
At each client, there consists of a frozen LLM model $f({\bm{x}};{\bm{\theta}})$ with a global lightweight global adapter (LoRA) $\Delta{\bm{\theta}}_{g}=\Delta{\bm{\theta}}^{b}_{g}\Delta{\bm{\theta}}^{a}_{g}$. This global adapter is used for aggregation by sending to the server. Notably, the server’s role is limited to computing the aggregated adapter $\Delta{\bm{\theta}}_{g}$, thus obviating the need for maintaining a large-scale model. Similar to the standard FL process, for each client $m$, the adapter weight $\Delta{\bm{\theta}}_{g}^{m}$ is learned locally and sent to the server. Upon receipt of the adapter weights from all clients, the server employs FedAvg mcmahan2017communication to aggregate them and sends $\Delta\bar{{\bm{\theta}}_{g}}$ back to each client as their initialized parameter in a new round. It can be formulated as:
[…]
@@
Other federated algorithms like FedProx li2020federated can also be applied with LoRA tuning of this global model learning (in Appendix B.1). In this paper, we just take FedAvg as an example.
@@ 4.2 Personalization of Local Model
The previously developed global model, which focuses on acquiring generic features across diverse datasets, faces challenges with personalization due to inconsistent optimization objectives. To address this, we integrate a local adapter to better align with each client’s specific distribution. We explore two methods as shown in Fig 2, 1) Learning sequentially: after global adapter training, the local adapter is initialized by the learned global adapter and directly fine-tuned; 2) Learning iteratively: during each communication round of global adapter training, the local adapter is re-initialized from its last state, fine-tuned alongside the frozen global adapter, and maintained locally without communication.
@@
Figure 2: Frameworks of two personalized methods for local adapter (LoRA) are shown on the left, with their overall learning processes on the right.
@@
To be more specific, a local adapter (LoRA) $\Delta{\bm{\theta}}_{l}=\Delta{\bm{\theta}}^{b}_{l}\Delta{\bm{\theta}}^{a}_{l}$ is introduced. Thus, each client contains three components: a frozen LLM ${\bm{\theta}}$, a global adapter (LoRA) $\Delta{\bm{\theta}}_{g}$ and a local adapter (LoRA) $\Delta{\bm{\theta}}^{m}_{l}$. As delineated in Fig 2 (a), for the first method, after global training, the local adapter is first initialized by the global adapter denoted as $\Delta{\bm{\theta}}^{m}_{l}=\Delta{\bm{\theta}}_{g}$, then fine-tuned on local data to get the final local adapter. As shown in Fig 2 (b), for the second method, during each communication round in training for each adapter layer, upon receiving an input ${\bm{h}}$, it simultaneously traverses the frozen LLM, the frozen global adapter and the local adapter. The process entails an initial fusion of the outputs from both the local and global adapters with a predefined weighting factor of $\alpha$, followed by integration with the output of the LLM to yield the final result ${\bm{h}}^{{}^{\prime}}={\bm{\theta}}{\bm{h}}+((1-\alpha)\cdot\Delta{\bm{\theta}}_{g}{\bm{h}}+\alpha\cdot\Delta{\bm{\theta}}^{m}_{l}{\bm{h}})$. Therefore, the learning of the local adapter $\Delta{\bm{\theta}}^{m}_{l}$ for these two methods can be unified as:
$\Delta{\bm{\theta}}^{m}_{l}=\text{arg}\min_{\Delta{\bm{\theta}}^{m}_{l}}\mathcal{L}_{m}({\bm{\theta}};\Delta{\bm{\theta}}_{g};\Delta{\bm{\theta}}^{m}_{l}),\quad\text{initialized with $\Delta{\bm{\theta}}_{g}$ or previous $\Delta{\bm{\theta}}_{l}^{m}$}$ (3)
4.3 LLM-enhanced Instance-wise Dynamic Weighting Mechanism
@@
As discussed in previous test-time FL methods jiang2023test , a dynamic combination of global components and personalized components can improve generalization while reducing the cost of hyper-parameter tuning in the deployment stage. Considering the disparate data distributions that characterize test-time tasks and local tasks and the wealth of training instances of local tasks available to each client, we propose an instance-wise dynamic weighting mechanism to calculate the similarity between the input instance and local instances, using this metric to determine the appropriate weight balance for the global and local adapter combination. To facilitate this, the representation of each input instance is essential. Leveraging the robust capability of pre-trained LLMs to abstract input sentences, we utilize the hidden states from the final layer of the LLM as the representation. Given that the LLM is decoder-based, with tokens attending only to preceding tokens, the embedding of the final token is considered representative of the entire input for similarity evaluation. Furthermore, to enhance the representation quality, the global adapter, which embodies generic knowledge, is incorporated into this embedding process.
@@
More specifically, during the inference stage, for each input instance ${\bm{x}}$ in a client, we randomly sample $S$ instances $\{{\bm{x}}_{0},{\bm{x}}_{1},...,{\bm{x}}_{s}\}$ from the local training dataset. These instances are then fed into the LLM, augmented with the global adapter, to obtain the last token’s embeddings from the final layer, denoted as ${\bm{w}}_{x}$ and $\{{\bm{w}}_{x_{0}},{\bm{w}}_{x_{1}},...,{\bm{w}}_{x_{s}}\}$ respectively. Subsequently, we calculate the similarity between the input representation ${\bm{w}}_{x}$ and each sampled local representation in $\{{\bm{w}}_{x_{0}},{\bm{w}}_{x_{1}},...,{\bm{w}}_{x_{s}}\}$, resulting in a score range of $[0,1]$. Finally, we average all scores to obtain the final result, represented as $\alpha_{t}=\lambda\cdot\sum_{i=0}^{S}\frac{1}{S}\operatorname*{Sim}({\bm{w}}_{x},{\bm{w}}_{x_{i}})$, where $\operatorname*{Sim}$ represents the function to calculate the similarity, and $\lambda$ is a scale factor in $(0,1]$ to restrict the maximum similarity score (especially for FedDPA-T).
@@
Through this method, the balancing of weights between the global and local adapters is dynamically adjusted for each test instance, ensuring the model not only tailors to the individual client’s specific needs but also benefits from the aggregated model’s generic knowledge across test-time tasks.
5 Experiment
5.1 Experiment Setting
Datasets.
We construct two federated datasets from Flan wei2021finetuned , which is a collection of various NLP tasks from over 60 datasets for instruction turning. In order to be better suitable for FL settings, we randomly select 8 NLP tasks from different datasets for each federated dataset and downsample the original datasets with more details in Appendix A.1. ROGUE-1 is taken as a metric.
Baselines and Implementation.
We compare our methods with four baselines based on the same model architecture: centralized model, Local-finetuned model, FedIT zhang2023towards and FedLoRA yi2023fedlora . The centralized model is trained on all data of tasks in one center. The local-finetuned model infers that only local data are used to train the model without any communication with other clients or the server.
We distribute data between clients based on the NLP task for data heterogeneity, where different NLP tasks generated from different contextual factors inherently suffer from various complex distribution shifts. Since we select 8 NLP tasks, corresponding to $M=8$ clients.For each client, the local task serves as the primary focus for personalization, while the tasks from other clients are taken as test-time tasks. To better evaluate the effectiveness of methods, we assume that all clients are activated for every communication round and set the communication round $K=20$. The alpaca-LoRA [^2] https://github.com/tloen/alpaca-lora is adapted as the base model initialized with LLaMA-7B. [^3] https://huggingface.co/huggyllama/llama-7b The updating weight of local LoRA training (FedDPA-T) is $\alpha=0.5$ ($\lambda=0.5$) for federated dataset 1 and $\alpha=0.3$ ($\lambda=0.3$) for federated dataset 2. We set $S=5$ and choose cosine similarity for instance-wise dynamic weighting mechanism. More details are in Appendix A.2.
Table 1: Personalization and test-time personalization results of different models on federated dataset 1. FedDPA-F represents the model with the local fine-tuning adapter and FedDPA-T represents the model with the local training adapter. Linguistic represents the linguistic acceptability task, Word Dis represents the word disambiguation task, and Question CLS represents question classification task.
@@ 5.2 Main Results
We compare FedDPA with other baselines on two main evaluation facets: personalization (scores on targeted local tasks) and test-time personalization (average scores on all tasks including test-time tasks). As evidenced in Table 1 and Table 2, our proposed dual-personalizing adapter methods (both fine-tuning and training) exhibit superior performance in personalization compared to other baseline models, which demonstrates the effectiveness of local adapter maintenance for enhancing performance on the targeted local task. For test-time personalization, the FedDPA-F method stands out as the most effective among all personalized models, which suggests that incorporating learning from the global adapter can be instrumental in adapting to test-time distribution shifts for a more comprehensive model achievement. Additionally, given that the global adapter aggregated on different distributions matin certain generalization capabilities, the local adapter of FedDPA-F has better generalization performance than that of FedDPA-T, which leads to better performance on most test-time tasks. More importantly, it is noteworthy that while centralized or global models may yield higher average performances across all tasks, they fall short in excelling at specific tasks for personalization, aligning with the conclusions of the previous study wang2023far .
@@ 6.1 Convergence Analysis
We present the convergence analysis of our methods in Figure 6.1. Figure 6.1 (a) compares our methods with other baselines for personalization, with the results showcasing the average performance on target local tasks across all clients. Notably, our methods exhibit a more rapid convergence compared to FedIT and achieve notable performance enhancements after five communication rounds. Despite sharing similar trends with FedLoRA, our approaches, particularly the FedDPA-T, ultimately outperform in personalization. For a more granular insight into test-time personalization convergence, Figure 6.1 (b) compares average performance on all tasks, including each client’s targeted local and test-time tasks. The results substantiate that our approaches demonstrate faster convergence rates, further bolstering the efficacy of our methods.
@@
Figure 3: Average accuracy varies as communication rounds.
Figure 4: Average accuracy varies as different client participation numbers.
6.2 Ablation Study
Impact of Instance-Wise Dynamic Weighting Mechanism.
To explore the impact of the instance-wise dynamic weighting mechanism, we implemented experiments with FedDPA methods on different datasets. As shown in Table 6.1, the incorporation of an instance-wise dynamic weighting mechanism contributes significantly to enhancing performance in both personalization and test-time personalization scenarios. More ablation studies are in Appendix B.2.
Impact of Updating Weight $\alpha$.
In this study, we investigated the influence of the updating weight $\alpha$ during FedDPA-T training with its value $\alpha\in\{03,0.5,0.7\}$. As can be seen in Table 6.1, for test-time personalization, increasing updating weight $\alpha$ will decrease the performance due to the increased proportion of the local adapter in the model, while for personalization, different updating weights $\alpha$ are required for different datasets to achieve their optimal results.
Impact of Client Number.
To better align with the FL setting in practical application, we scaled up clients to 40 and implemented experiments with sample rate $\{0.2,0.4,0.6,0.8,1\}$. For each communication round, the server will select clients from each task based on the sample rate (more details in Appendix A.1). As shown in Figure 6.1, as the client participant rates increase, model accuracy also increases as more participating clients provide more data for knowledge learning. Besides, FedDPA-F outperforms all baselines, whereas FedDPA-T exhibits somewhat inferior performance, potentially due to overfitting issues when handling a small dataset.
More experiments and analyses of scalability and efficiency can be found in Appendix B.3 and B.4.